OpenStudy (anonymous):

Use the square root 2x^(2)-14=0

OpenStudy (anonymous):

you want to complete the square, you mean?

OpenStudy (anonymous):

and solve f or x

OpenStudy (anonymous):

2x^2=14 x^2=7 x=\[\sqrt{7}\]

OpenStudy (anonymous):

resolves?

OpenStudy (mertsj):

positive or negative sqrt7

OpenStudy (anonymous):

2x^(2) = 14 (2x^(2))^(1/2) = +/- (14)^(1/2) x = +/- (14)^(1/2)/2

OpenStudy (anonymous):

Mertsj +or- :)

OpenStudy (anonymous):

fontez it is a squre so it would be +/-

OpenStudy (anonymous):

This is what I got. √(7), -√(7)

OpenStudy (anonymous):

yep leidyr

OpenStudy (anonymous):

yeah you are right I could have simplified it further

OpenStudy (anonymous):

remember if you take the cubic, or odd root you only get one answer

OpenStudy (anonymous):

decimals numbers :: x=2,6... or x=-2,6... :)